OpenStudy (anonymous):

@AdrDiazEs Indicate in standard form the equation of the line through the given points. K(6, 4), L(-6, 4) What formula do I use?

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

this is what a standard form looks like, \[\huge ax+by = c\]

OpenStudy (anonymous):

first find the slope of a line. do you know how to find the slope ?

OpenStudy (anonymous):

ermm, I probably have before but I dont remember. /.\

OpenStudy (anonymous):

wait wait lemme check

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\] ?

OpenStudy (anonymous):

yes, you're on the right track, now plug the points and solve to find the slope

OpenStudy (anonymous):

I think I messed up lol... I got \[\frac{ 4 - 4 }{ -6 - 6 }\]

OpenStudy (anonymous):

which is 0/-12

OpenStudy (anonymous):

yes and \[\frac{ 0 }{ -12 } = 0 \] so the slope would be 0

OpenStudy (anonymous):

now you have the slope, use a tangent line formula of \[\huge y-y_1 = m(x-x_1)\] m is the slope and plug in the points

OpenStudy (anonymous):

\[\huge (6,4) = (\color {red}{x_1, y_1 })\]

OpenStudy (anonymous):

( 6 , 4 ) = (6,4) ?

OpenStudy (anonymous):

i meant to say, you can use those points and plug them into the tangents line formula

OpenStudy (anonymous):

ooohhh lol. y-4=0(x-6) ?

OpenStudy (anonymous):

now distribute the \[\huge 0(\color{green}{x-6}) \]

OpenStudy (anonymous):

the zero to x-6

OpenStudy (anonymous):

so you would have \[\huge y-4 = 0\]

OpenStudy (anonymous):

now try to put that in to the standard form that i gave you at the beginning

OpenStudy (anonymous):

x + y - 4 + 0 ?

OpenStudy (anonymous):

in this case the